A sinusoidal voltage of 120 V is applied to a resistor of 10 Ohm.
a) calculate the effective current in the registor
b) the peak voltage across the registor
Answer #1

a.) The effective current is given as V/R

Here V = 120 volts and R = 10 ohms

Therefore, current = 120/10 = 12 amperes.

b.) Assuming the given resistor is the only resistor in the circuit, the maximum voltage across it will be equal to the maximum voltage applied across the circuit, which in this case is 120 Volts.


answered by: ANURANJAN SARSAM
Answer #2

a.) The effective current is given as V/R

Here V = 120 volts and R = 10 ohms

Therefore, current = 120/10 = 12 amperes.

b.) Assuming the given resistor is the only resistor in the circuit, the maximum voltage across it will be equal to the maximum voltage applied across the circuit, which in this case is 120 Volts.